A regular octagon of side 10 is inscribed in a circle. Find the radius of the circle.


1
Expert's answer
2021-08-24T15:51:21-0400
"\alpha=360\degree/8=45\degree"

The Law of Cosines


"10^2=r^2+r^2-2r^2\cos45\degree"

"r^2=\dfrac{100}{2-\sqrt{2}}"

"r=10\sqrt{1+\dfrac{\sqrt{2}}{2}}"